OpenStudy (fibonaccichick666):

How do you apply vieta's formula to \[1+z+...+z^9=0?~ z^{10}=1~ z\not=1\]

OpenStudy (anonymous):

\[1+z+z^2+...+z^{10}=1\] making one solution \(z=0\) and the other the tenth roots of 1 i think

OpenStudy (fibonaccichick666):

well, the prob is i have (z-1)(1+z+...z^9)=z^10-1

OpenStudy (anonymous):

i guess this is not viete's formula

OpenStudy (fibonaccichick666):

but I don't really understand the theorem

OpenStudy (fibonaccichick666):

http://en.wikipedia.org/wiki/Vieta's_formulas

OpenStudy (fibonaccichick666):

I mean I can make that relation

OpenStudy (anonymous):

not sure what you are supposed to do with them, they are the elementary symmetric functions it does not solve the equation for you though

OpenStudy (fibonaccichick666):

ugh, grrr I'm hung up on the pf, for tan pi/10

OpenStudy (fibonaccichick666):

i wanted to use the complex way, but I don't understand how they went from step to step in the pfs you supplied

OpenStudy (anonymous):

did you get any help for zakon?

OpenStudy (fibonaccichick666):

nope, he said you were rigth

OpenStudy (anonymous):

*from

OpenStudy (anonymous):

lol great

OpenStudy (fibonaccichick666):

yup haha

OpenStudy (fibonaccichick666):

do you think you could explain the pf? or at least help with the steps, I just don't get it

OpenStudy (anonymous):

ok lets see if we can figure it out

OpenStudy (fibonaccichick666):

it's this If 10x=π sin2x=cos3x as 2x+3x=5x=π2 ⟹2sinxcosx=4cos3x−3cosx ⟹2sinx=4cos2x−3 as cosx≠0 If sinx=t,2t=4(1−t2)−3⟹4t2+2t−1=0 ⟹t=−1±5‾‾√4 , but sinx>0 as 0<x<π sinπ10=5‾‾√−14 (1)So, cosπ10=1−(sinπ10)2‾‾‾‾‾‾‾‾‾‾‾‾‾√=10+25‾‾√‾‾‾‾‾‾‾‾‾‾√4 So, tanπ10=5√−1410+25√√4=5‾‾√−110+25‾‾√‾‾‾‾‾‾‾‾‾‾√=5‾‾√−110+25‾‾√‾‾‾‾‾‾‾‾‾‾√ =(5‾‾√−1)210+25‾‾√‾‾‾‾‾‾‾‾‾‾‾√=3−5‾‾√5‾‾√(5‾‾√+1)‾‾‾‾‾‾‾‾‾‾‾‾‾√=(3−5‾‾√)(5‾‾√−1)5‾‾√(5‾‾√+1)(5‾‾√−1)‾‾‾‾‾‾‾‾‾‾‾‾‾‾‾‾‾‾‾‾‾√=5‾‾√−25‾‾√‾‾‾‾‾‾‾‾√ Or(2) cosπ5=1−2(5‾‾√−14)2=5‾‾√+14 We know cos2y=1−tan2y1+tan2y⟹tan2y=1−cos2y1+cos2y So, tan2π10=1−5√+141+5√+14=3−5‾‾√5‾‾√(5‾‾√+1) which we have already encountered in (1).

OpenStudy (fibonaccichick666):

so first question, how did he do the 10x=pi

OpenStudy (anonymous):

ok hold the phone, maybe there is an easier way

OpenStudy (fibonaccichick666):

ok i'll wait

OpenStudy (fibonaccichick666):

ooh nvm! got it!

OpenStudy (fibonaccichick666):

Thanks sat!

OpenStudy (anonymous):

oh, maybe this is snappy enough to do it the other way, you make \(\sin(5x)\) in to a quadratic equation in \(\sin(x)\) oh ok good because although this is a gimmick it is pretty snappy right?

OpenStudy (fibonaccichick666):

it is it is haha

OpenStudy (fibonaccichick666):

young snappy snapper

OpenStudy (anonymous):

lol